The eye of a hurricane passes over Grand Bahama Island. It is moving in a direction 62.3◦ north of west with a speed of 45.5 km/h. Exactly 3.07 hours later, the course of the hurricane shifts due north, and its speed slows to 29.1 km/h, as shown. How far from Grand Bahama is the hurricane 5.31 h after it passes over the island?
1
Expert's answer
2012-10-19T08:48:28-0400
L(west) = 3.07 * 45.5 * cos(62.3◦) = 64.93 km
L(north) = 3.07 * 45.5 *
sin(62.3◦) + (5.31 - 3.07) * 29.1 = 123.68 +
65.184 = 188.86 km
D = sqrt(
L(west)^2 + L(north)^2 ) = sqrt( 64.93^2 + 188.86^2 ) = 200 km